Saving for Dream Vacation. Kimberly makes $8,000 more a year than her husband. Kimberly saves 10% of her salary for vacation, her husband saves 8%. Together the saved $5,660 per year. How much does Kimberly and her husband make?
Answer by asha(30)

let the husband earn 100x$.Kimberley earns 100x+8000$
husband saves8% of his salary =8x$
Kimberly saves 10% of her salary =1/10(100x+8000)=10x+800
Together they save 8x+10x+800 =18x +800 =5660
18x=5660-800=4860
x=4860/18=270
Kimberly earns 270*100+8000=27,000+8000=35000$
Her husband earns 27000$